What percent of the area under the normal curve is between z=2.25 and z=0.85

1
Expert's answer
2022-02-28T13:52:06-0500

"P(0.85<Z<2.25)=P(Z<2.25)-P(Z<0.85)=0.9878-0.8023=0.1854."